A train is moving with a velocity of 30 m/s. When brakes are applied, it is found that the velocity reduces to 10 m/s in 240 m. When the velocity of the train is zero, the distance travelled is

Grade:11

1 Answers

Vikas nayak
2 years ago
Now, let us take Final velocity and initial velocity 
V^2- U^2= 2as  {v=10, u=30} 
a= -800/2*240 = -5/3 (v=0), 
Now again apply equation, 
V^2-U^2= 2as
0-u^2=2*-5/3*S 
S= 900*3/10 = 270m
So total distance travelled is 270m ..